# Polyhedral Complex Derivation from Piecewise Trilinear Networks

### v1

@@ 4.2 Curved Edge Subdivision in a Trilinear Space
$\{{\bm{x}}:\tau({\bm{x}})=(1-t)\tau({\bm{x}}_{0})+t\tau({\bm{x}}_{7})\hskip 40.00006pt$
[…]
@@
The detailed proof is provided in Lemma A.4. ∎
@@
A hypersurface $\tau({\bm{x}})=0$ passing two points $\tau({\bm{x}}_{0})=\tau({\bm{x}}_{7})=0$ while $\tau({\bm{x}}_{1\dots 6})\neq 0$ satisfies the eikonal constraint $\|\nabla\tau({\bm{x}})\|_{2}=1$ for all ${\bm{x}}\in[0,1]^{3}$. Then, the hypersurface of $\tau({\bm{x}})=0$ is a plane.
@@
The proofs can be found in Theorem A.5 and Corollary A.6
Since we aim for a polyhedral complex derivation that piecewise trilinear networks represent, we replace one of the hypersurfaces (possibly) forming the curved edge with a diagonal plane in a piecewise trilinear region. We opt for this choice not only due to its mathematical simplicity but also the characteristics of trilinear hypersurfaces as illustrated in Figure 6, Appendix. Thus, the new vertices lie on at least two hypersurfaces, and the new edges exist on the same hypersurface, while the eikonal constraint minimizes any associated error. This error is tolerated by the hyperparameter $\epsilon$ as specified in Definition 3.4.
@@
Let $\tilde{\nu}_{0}({\bm{x}})=0$ and $\tilde{\nu}_{1}({\bm{x}})=0$ be two hypersurfaces passing two points ${\bm{x}}_{0}$ and ${\bm{x}}_{7}$ such that $\tilde{\nu}_{0}({\bm{x}}_{0})=\tilde{\nu}_{1}({\bm{x}}_{0})=0$ and $\tilde{\nu}_{0}({\bm{x}}_{7})=\tilde{\nu}_{1}({\bm{x}}_{7})=0$, ${\bm{P}}_{i}:=\tilde{\nu}_{0}({\bm{x}}_{i})$ and ${\bm{Q}}_{i}:=\tilde{\nu}_{1}({\bm{x}}_{i})$, ${\bm{P}}_{\alpha}=\big{[}{\bm{P}}_{0};~{}{\bm{P}}_{1};~{}{\bm{P}}_{4};~{}{\bm{P}}_{5}\big{]}$, ${\bm{P}}_{\beta}=\big{[}{\bm{P}}_{2};~{}{\bm{P}}_{3};~{}{\bm{P}}_{6};~{}{\bm{P}}_{7}\big{]}$, and ${\bm{X}}=[(1-x)^{2};~{}x(1-x);~{}(1-x)x;~{}x^{2}]$. Then, $x,z\in[0,1]$ of the intersection point of the two hypersurfaces and a diagonal plane of $x=z$ is the solution of the following quartic equation:
@@
Please refer to Theorem A.7 for the proof. ∎
@@
Note that finding roots of a polynomial is the eigenvalue decomposition of a companion matrix (Edelman & Murakami, 1995), which can be parallelized (ref. Lemma A.8). Given that the companion matrix remains small for a quartic equation in $\mathbb{R}^{4\times 4}$, the overall complexity remains $\mathcal{O}(|\mathcal{V}|)$. For the detailed complexity analysis on the proposed method, please refer to Appendix B.
4.3 Skeletonization
Skeletonization selects the vertices such that:
${\mathcal{V}}^{\star}=\{{\bm{x}}|\tilde{\nu}({\bm{x}})\leq\epsilon,{\bm{x}}\in{\mathcal{V}}\}$ (18)
@@ 4.3 Skeletonization
and the edges ${\mathcal{E}}^{\star}$ such that their two vertices are among ${\mathcal{V}}^{\star}$.
4.4 Faces and Normals
@@
Edges lying on the same plane and a shared region form a face, and this alignment can be determined by evaluating the sign-vectors (ref. Definition 3.4 and Section 5.2). It is important to note that faces are not inherently triangular; however, if needed, they can be triangulated via triangularization for further analysis. The order of vertices conventionally describes its normal. Let ${\bm{x}}_{0}$, ${\bm{x}}_{1}$, and ${\bm{x}}_{2}$ be the vertices forming a triangular face. The normal is defined as follows:
${\bm{n}}=({\bm{x}}_{1}-{\bm{x}}_{0})\times({\bm{x}}_{2}-{\bm{x}}_{0})$ (19)
where $\times$ denotes cross product, and the normal is orthogonal to both vectors, with a direction given by the right-hand rule. If the viewing direction aligns opposite to the normal, the order of vertices follows a counter-clockwise arrangement.
We determine the vertex order based on the normal derived from the Jacobian of the mean of vertices. Using cosine similarity to compute relative angles $\theta\in[-\pi,\pi]$ and cross product for getting direction, we identify the counter-clockwise arrangement of vertices, as outlined in Algorithm 1.

### v2

@@ 4.2 Curved edge subdivision in a trilinear space
$\{{\bm{x}}:\tau({\bm{x}})=(1-t)\tau({\bm{x}}_{0})+t\tau({\bm{x}}_{7})~{}~{}\text{where}~{}~{}{\bm{x}}\in\mathbb{R}^{D}~{}~{}\text{and}~{}~{}t\in\mathbb{R}\}.$
@@
The detailed proof using the piecewise linearity of $\nu$ is provided in Lemma D.4. ∎
@@
A hypersurface $\tau({\bm{x}})=0$ intersects two points $\tau({\bm{x}}_{0})=\tau({\bm{x}}_{7})=0$ while $\tau({\bm{x}}_{1\dots 6})\neq 0$ for the remaining six points. These points form a cube, with ${\bm{x}}_{0}$ and ${\bm{x}}_{7}$ positioned on the diagonal of the cube. The hypersurface satisfies the eikonal constraint $\|\nabla\tau({\bm{x}})\|_{2}=1$ for all ${\bm{x}}\in[0,1]^{3}$. Then, the hypersurface of $\tau({\bm{x}})=0$ is a plane.
@@
The proofs using its partial derivatives can be found in Theorem D.5 and Corollary D.6.
Since we aim for a practical polyhedral complex derivation that piecewise (still) trilinear networks represent, we replace one of the hypersurfaces forming the curved edge with a diagonal plane in a piecewise trilinear region. We opt for this choice not only due to its mathematical simplicity but also the characteristics of trilinear hypersurfaces as illustrated in Figure 4, Appendix G. Thus, the new vertices lie on at least two hypersurfaces, and the new edges exist on the same hypersurface, while the eikonal constraint minimizes any associated error from this approximation. This error is tolerated by the hyperparameter ϵ=italic-$\epsilon=$1e-4 as specified in Definition 3.4 and Section 6. More discussions on this matter can be found in Section B.1.
@@
Let $\tilde{\nu}_{0}({\bm{x}})=0$ and $\tilde{\nu}_{1}({\bm{x}})=0$ be two hypersurfaces passing two points ${\bm{x}}_{0}$ and ${\bm{x}}_{7}$ such that $\tilde{\nu}_{0}({\bm{x}}_{0})=\tilde{\nu}_{1}({\bm{x}}_{0})=0$ and $\tilde{\nu}_{0}({\bm{x}}_{7})=\tilde{\nu}_{1}({\bm{x}}_{7})=0$, ${\bm{P}}_{i}:=\tilde{\nu}_{0}({\bm{x}}_{i})$ and ${\bm{Q}}_{i}:=\tilde{\nu}_{1}({\bm{x}}_{i})$, ${\bm{P}}_{\alpha}=\big{[}{\bm{P}}_{0};~{}{\bm{P}}_{1};~{}{\bm{P}}_{4};~{}{\bm{P}}_{5}\big{]}$, ${\bm{P}}_{\beta}=\big{[}{\bm{P}}_{2};~{}{\bm{P}}_{3};~{}{\bm{P}}_{6};~{}{\bm{P}}_{7}\big{]}$, and ${\bm{X}}=[(1-x)^{2};~{}x(1-x);~{}(1-x)x;~{}x^{2}]$. Then, $x\in[0,1]$ of the intersection point of the two hypersurfaces and a diagonal plane of $x=z$ is the solution of the following quartic equation:
@@
Please refer to Theorem D.7 for the detailed proof, which uses linear algebra to rearrange two trilinear equations with two variables to get a solution. ∎
@@
Note that finding roots of a polynomial is the eigenvalue decomposition of a companion matrix [29], which can be parallelized (ref. Lemma D.8). Given that the companion matrix remains small for a quartic equation in $\mathbb{R}^{4\times 4}$, the overall complexity remains $\mathcal{O}(|\mathcal{V}|)$. For the detailed complexity analysis on the proposed method, please refer to Appendix E.
4.3 Skeletonization, faces, and normals
@@
Skeletonization selects the vertices such that: ${\mathcal{V}}^{\star}=\{{\bm{x}}|\tilde{\nu}({\bm{x}})\leq\epsilon,{\bm{x}}\in{\mathcal{V}}\}$, and the edges ${\mathcal{E}}^{\star}$ such that their two vertices are among ${\mathcal{V}}^{\star}$.
@@
Edges lying on the same plane and a shared region form a face, and this alignment can be determined by evaluating the sign-vectors (ref. Definition 3.4 and Section 5). Note that faces are not inherently triangular; however, if needed, they can be triangulated via triangularization for further analysis. The order of vertices conventionally describes its normal. Let ${\bm{x}}_{0}$, ${\bm{x}}_{1}$, and ${\bm{x}}_{2}$ be the vertices forming a triangular face. The normal is defined using cross-product as ${\bm{n}}=({\bm{x}}_{1}-{\bm{x}}_{0})\times({\bm{x}}_{2}-{\bm{x}}_{0})$ where the normal is orthogonal to both vectors, with a direction given by the right-hand rule. If the viewing direction aligns opposite to the normal, the order of vertices follows a counter-clockwise arrangement. We determine the vertex order based on the normal derived from the Jacobian of the mean of vertices. Using cosine similarity to compute relative angles $\theta\in[-\pi,\pi]$ and cross product for getting direction, we identify the counter-clockwise arrangement of vertices, as outlined in Algorithm 3 in Appendix.
5 Implementation
We employ the HashGrid [12] for $\tau$, although any encoding using trilinear interpolation we defined can be applied, while our discussion can be found in Section B.2. We illustrate our method in Algorithm 1 and 2, describing the details in the following sections.
Initialization.
While the original edge subdivision algorithm [4] starts with a unit cube of eight vertices and a dozen edges, we know that the unit cube is subdivided by orthogonal planes consisting of the grid. Let an input coordinate be ${\bm{x}}\in[0,1]^{3}$, and there is a unit cube such that one of its corners is at the origin. Taking into account multi-resolution grids, we obtain the marks $m_{i}$, such that the grid planes are $x=m_{i}$, $y=m_{i}$, and $z=m_{i}$, following Algorithm 4 in Appendix. Notice that we carefully consider the offset $s/2$, which prevents the zero derivatives from aligning across all resolution levels (ref. Appendix A of Müller et al. [12]). Leveraging the obtained marks, we derive the associated grid vertices and edges, which serve as the initial sets for ${\mathcal{V}}$ and ${\mathcal{E}}$, respectively.
Optimizing sign-vectors.
Piecewise trilinear region.
In Theorem 4.7, we need the outputs of corners ${\bm{P}}$ and ${\bm{Q}}$ in a common piecewise trilinear region; however, some corners may be outside of the region. For this, we replace all ReLU activations using the mask ${\bm{m}}$ that: ${\bm{m}}_{j}^{(i)}=(\tilde{\nu}_{j}^{(i)}({\bm{x}}_{0})>\epsilon)~{}|~{}(\tilde{\nu}_{j}^{(i)}({\bm{x}}_{7})>\epsilon),$ where ${\bm{x}}_{0}$ and ${\bm{x}}_{7}$ are the vertices of an edge of interest, for the calculated cubic corners ${\bm{x}}_{0\dots 7}$. This implies that $\nu$ exhibits linearity, except for instances where two vertices simultaneously deviate from linearity. For ${\bm{P}}$ and ${\bm{Q}}$, we select the last two pairs of $i$ and $j$ such that $\gamma_{\phi(i,j)}({\bm{x}}_{0})=\gamma_{\phi(i,j)}({\bm{x}}_{7})=0$, indicating two common hypersurfaces passing two points ${\bm{x}}_{0}$ and ${\bm{x}}_{7}$.
Algorithm 1 Polyhedral Complex Derivation
  Input: Sets of vertices ${\mathcal{V}}$ and edges ${\mathcal{E}}$ (from Section 5), # of layer $L$, hidden size $H$, trilinear networks $\tilde{\nu}$
  for $i=0$ to $L-1$ do
     for $j=0$ to $H-1$ do
        ${\mathcal{V}},{\mathcal{E}}=$ CESub$({\mathcal{V}},{\mathcal{E}},\tilde{\nu},i,j)$triangleright Algorithm 2
     end for
  end for
  ${\mathcal{V}},{\mathcal{E}}=$ CESub$({\mathcal{V}},{\mathcal{E}},\tilde{\nu},L,0)$
  return ${\mathcal{V}}^{\star},{\mathcal{E}}^{\star},{\mathcal{F}}^{\star}$
Algorithm 2 Curved Edge Subdivision (CESub)
  Input: Vertices ${\mathcal{V}}$, Edges ${\mathcal{E}}$, trilinear networks $\tilde{\nu}$, layer index $i$, neuron index $j$
  $d=\tilde{\nu}({\mathcal{V}};i,j)$, ${\mathcal{V}}_{\text{new}}=\emptyset$
  for $e$ in ${\mathcal{E}}$ do
     if $d_{e_{0}}\cdot d_{e_{1}}<0$ then
        ${\bm{P}},{\bm{Q}}=$ Corners$(e)$triangleright Section 5
        ${\bm{x}}_{\text{new}}=$ Intersect$({\bm{P}},{\bm{Q}},e)$triangleright Theorem 4.7
        $e\leftarrow[e_{0},~{}\text{Index}({\bm{x}}_{\text{new}})]$
     end if
  end for
  ${\mathcal{V}}\leftarrow{\mathcal{V}}+{\mathcal{V}}_{\text{new}}$
  return ${\mathcal{V}}$, ${\mathcal{E}}$
6 Experiment
Objective.
For a given trilinear neural network with the eikonal constraint, our goal is to extract meshes from it in a way that accurately reflects the learned SDF. (Note that NeuS [6] showed how to convert the density networks of NeRFs to an SDF in the frameworks of volume rendering.) To evaluate this, we utilize marching cubes with excessive sampling to get a pseudo-ground truth mesh.
Hyperparameters.
For ReLU neural networks, the number of layers $L$ of 3, hidden size $H$ of 16, and $\epsilon$ of 1e-4 for the sign-vectors (ref. Definition 3.4). The weight for the eikonal loss is 1e-2. For HashGrid, the resolution levels of 4, feature size of 2, base resolution $N_{\text{min}}$ of 2, and max resolution $N_{\text{max}}$ of 32 by default. $N_{\text{min}}$ and $N_{\text{max}}$ are doubled (x2) or quadrupled (x4) for Medium or Large settings in Figure 3. We use the official Python package of tinycudnn  [^4] https://github.com/NVlabs/tiny-cuda-nn for the HashGrid module.
Chamfer distance.
We randomly select 100K points on the faces through ray-marching from the origin, directing toward a randomly chosen point on a unit sphere.
Table 1 and Figure 3 show the results for the Standford 3D Scanning repository [30]. Varying the number of samples in marching cubes [9], we plot the baseline with respect to the number of generated vertices. Given that our method extracts vertices from the intersection points of hypersurfaces, it enables parsimonious representations using the lower number of vertices having better CD, summarized by chamfer efficiency (CE) in Table 1, and ours are located in the lower-left region compared with the baselines in Figure 3. In particular, we observe a strong CE in a simple object, e.g., Drill Bit, which achieved a CE of 47.8 versus 19.2 (MC).
Our method inherently holds an advantage over MC in capturing optimal vertices and faces, especially in cases where the target surfaces exhibit planar characteristics. When dealing with curved surfaces, the smaller grid size of the Large model in Figure 3 views a curved surface in a smaller interval, and it tends to approach the plane. From the plotting of Small to Large, our CDs (colored crosses) are decreasing to zero, consistently with better CEs, confirming this speculation. The complete results of the Large models for the Stanford 3D Scanning repository can be found in Tables 6, 7 and 8 in Appendix F and its visualization can be found in the right side of Figure 5.
Table 1: Chamfer distance (CD) and chamfer efficiency (CE) for the Stanford 3D Scanning repository [30]. The CD ($\times$1e-6) is evaluated using the marching cubes (MC) with 2563 grid samples as the reference ground truth for all measurements. The CE is defined as $100/(\sqrt{|{\mathcal{V}}|\times\text{CD}})$, providing a concise representation of the trade-off between the number of vertices and the CD. Please refer to Table 2 and Table 3 for the full results, and Table 4 for standard deviation and time spent in Appendix (Ours took 0.97 $\pm$ 0.21 while MC also took within 1 sec for the bunny.)
Bunny Dragon Happy Buddha Armadillo Drill Lucy
MC # $|{\mathcal{V}}|$downarrow CD $\downarrow$ CE $\uparrow$ $|{\mathcal{V}}|$downarrow CD $\downarrow$ CE $\uparrow$ $|{\mathcal{V}}|$downarrow CD $\downarrow$ CE $\uparrow$ CE $\uparrow$ CE $\uparrow$ CE $\uparrow$
  32   1283 4106 19.0   1416 6330 11.2   1032 6951 13.9 11.2   7.5 24.9
  64   5367 1371 13.6   6090 1936   8.5   4362 2465   9.3   8.3 19.2 16.0
128 21825   393 11.6 25236   542   7.3 18219   722   7.6   6.6 17.4 10.7
196 49569   141 14.3 57341   203   8.6 41351   276   8.8   7.5 15.3 11.7
Ours   4341   900 25.6   5104 1243 15.8   3710 1738 15.5 13.9 47.8 23.6
Figure 2: Chamfer distances for the Stanford dragon varying the number of vertices and the model sizes, showing consistent efficiency.
Figure 3: Effect of the weight of eikonal loss on the flatness error in Equation 13. For the plot, we conduct experiments using the unit sphere.
Angular distance.
The angular distance measures how much the normal vectors deviate from the ground truths. Analogous to the chamfer distance, we sample 100K points on the surface and calculate the normal vectors as described in Section 4.3. The angular distance is defined as follows:
In Table 5 in Appendix, the angular distances for the Stanford bunny are shown. As we expected, our approach efficiently estimates the faces using a parsimonious number of vertices compared to the sampling-based method. We confirm this trend is consistent across other objects.
Eikonal constraint for planarity.
To assess the efficacy of the eikonal constraint, which enforces the planarity of hypersurfaces as outlined in Theorem 4.5, we quantify the flatness error associated with the plenary constraints presented in the proof of Theorem D.5. Let ${\bm{x}}_{0}$ and ${\bm{x}}_{7}$ be two vertices consisting of a diagonal edge, while ${\bm{x}}_{1\dots 6}$ are its remaining cubic corners. We obtain $\tilde{\nu}({\bm{x}}_{1\dots 6})$ in a piecewise trilinear region as described in Section 5. This evaluation is conducted as follows:
In Figure 3, we empirically validate that the eikonal loss enforces planarity in the hypersurfaces within piecewise trilinear regions, as described in Theorem 4.5. The absence of the eikonal loss results in elevated planarity errors, leading to the construction of an inaccurate mesh.
Visualization.
We provide the qualitative analyses in Figures 5, 6, 7 and 8 in Appendix G encompassing the comparison between the Small and Large models (Figure 5), visualizing the marching cubes varying the number of samples (Figure 6), detailed comparisons (Figure 7), and the gallery of the Stanford 3D Scanning meshes from the Large models (Figure 8).
Limitations.
Although we implement the algorithm using pre-defined parallel tensor operations in PyTorch [31], achieving time and memory complexity of $\mathcal{O}(|\mathcal{V}|)$ as detailed in Appendix E, our method inherently requires forward passing the intermediate vertices during edge subdivisions iterating over neurons, which depends on previous steps (Algorithm 1). For the Stanford bunny using the Small model, the intermediate counts for vertices and edges exceed 11K and 19K, respectively, although they are eventually reduced to 4.3K vertices for face extraction. This overhead is negligible for the Small models, taking 0.97 seconds; however, for the Large models, the process takes 69.0 seconds for 136K vertices, compared to 11.88 seconds for marching cubes with 512 samples per axis. We note that parallelizing over multiple GPUs would be a reliable option to improve latency.
7 Conclusions
In conclusion, our exploration of novel techniques in visualizing trilinear neural networks unveils intricate structures and extends mesh extraction applications beyond CPWA functions. Focused on trilinear interpolating methods, our study yields novel theoretical insights and a practical methodology for precise mesh extraction. Notably, Theorem 4.5 shows the transformation of hypersurfaces to planes within the trilinear region under the eikonal constraint. Empirical validation using chamfer distance and efficiency, and angular distance affirms the correctness and parsimony of our methodology. The correlation analysis between eikonal loss and hypersurface planarity confirms the theoretical findings. Our proposed method for approximating intersecting points among three hypersurfaces, coupled with a vertex arrangement methodology, broadens the applications of our techniques. This work establishes a foundation for future research and practical applications in the evolving landscape of deep neural network analysis and toward real-time rendering mesh extraction.
Acknowledgments and Disclosure of Funding
I would like to express my sincere appreciation to my brilliant colleagues, Sangdoo Yun, Dongyoon Han, and Injae Kim, for their contributions to this work. Their constructive feedback and guidance have been instrumental in shaping the work. The NAVER Smart Machine Learning (NSML) platform [32] had been used for experiments.
Appendix A Tropical geometry and tropical algebra of neural networks
In this section, we present an overview of tropical geometry (Section A.1) to provide a formal definition of the tropical hypersurface (Section A.2). This hypersurface is significant as it serves as the decision boundary for ReLU neural networks optimizing an SDF, resulting in the formation of a polyhedral complex. For this, we delve into the examination of the tropical algebra of neural networks (Section A.3). Furthermore, we discuss the edge subdivision algorithm [4] for extracting the tropical hypersurface (Section 3.2), and we extend this discussion to include considerations for trilinear interpolation in Section 4. Given that Section A.1 to A.3 offer insights into tropical geometry, illustrating how each neuron contributes to forming the polyhedral complex, you may choose to skip these subsections if you are already acquainted with these concepts.
A.1 Tropical geometry
Tropical geometry [25, 26] describes polynomials and their geometric characteristics. As a skeletonized version of algebraic geometry, tropical geometry represents piecewise linear meshes using tropical semiring and tropical polynomial.
A.2 Tropical hypersurface
Definition A.1 (Tropical hypersurface, restated).
$V(f)=\{{\bm{x}}\in\mathbb{R}^{d}:c_{i}{\bm{x}}^{\mathrm{\alpha}_{i}}=c_{j}{\bm{x}}^{\mathrm{\alpha}_{j}}=f({\bm{x}})~{}~{}~{}~{}~{}\text{for some}~{}~{}~{}\mathrm{\alpha}_{i}\neq\mathrm{\alpha}_{j}\}$
where the tropical monomial $c_{i}{\bm{x}}^{\mathrm{\alpha}_{i}}$ with $d$-variate ${\bm{x}}$ is defined as $c\odot x_{1}^{a_{i,1}}\odot\cdots\odot x_{d}^{a_{i,d}}$ with $d$-variate $\mathrm{\alpha}_{i}=(a_{i,1},\dots,a_{i,d})$. In other words, this is the set of points where $f({\bm{x}})$ equals two or more monomials in $f$.
A.3 Tropical algebra of neural networks
Definition A.2 (ReLU neural networks, restated).
Although the neural networks $\nu$ is generally non-convex, it is known that the difference of two tropical signomials can represent $\nu$ [27]. A tropical signomial takes the same form as a tropical polynomial (ref. Definition 3.1), but it additionally allows the exponentials $\mathrm{\alpha}$ to be real values rather than integers. With tropical signomials $F$, $G$, and $H$,
$\nu^{(l)}({\bm{x}})$ $=H^{(l)}({\bm{x}})-G^{(l)}({\bm{x}})$ (15)
$\sigma^{(l)}\circ\nu^{(l)}({\bm{x}})$ $=F^{(l)}({\bm{x}})-G^{(l)}({\bm{x}})$ (16)
where $F^{(l)}$, $G^{(l)}$, and $H^{(l)}$ are as follows:
$F^{(l)}({\bm{x}})$ $=\max\big{(}H^{(l)}({\bm{x}}),~{}G^{(l)}({\bm{x}})\big{)}$ (17)
Proposition A.3.
(Decision boundary of neural networks, restated) The decision boundary for $L$-layer neural networks with ReLU activation $\nu^{(L)}({\bm{x}})=0$ is where two tropical monomials, or two arguments of $\max$, equal (Definition 3.1), recursively, as follows:
${\mathcal{B}}\subseteq\big{\{}{\bm{x}}:\nu_{j}^{(i)}({\bm{x}})=H_{j}^{(i)}({\bm{x}})-G_{j}^{(i)}({\bm{x}})=0,~{}i\in(1,\dots,L),~{}j\in(1,\dots,H)\big{\}}$
where the subscription $j$ denotes the $j$-th element, assuming the hidden size of neural networks is $H$, while the $L$-th layer has a single output for rather discussions as an SDF. Remind that tropical operations satisfy the usual laws of arithmetic for recursive consideration.
This implies that piecewise hyperplanes corresponding to the preactivations $\{\nu_{j}^{(i)}\}$ are the candidates shaping the decision boundary. Each linear region has the same pattern of zero masking by $\textrm{ReLU}(\cdot)$, effectively making it linear as an equation of a plane. In turn, if the pattern changes as an input is continually changed, it moves to another linear region. From a geometric point of view, this is why the number of linear regions grows polynomially with the hidden size $H$ and exponentially with the number of layers $L$ [27, 18].
Appendix B Further discussions
B.1 Satisfying the eikonal equation and approximation
Theorem 4.7 handles where the eikonal equation does not make perfect hyperplanes. First, we observed the characteristics of trilinear hypersurfaces as illustrated in Figure 4 in Appendix. Among 64 edge scenarios within a trilinear region, a diagonal plane intersects with the hypersurface in nearly all cases. Moreover, we can take advantage of it since we can eliminate one variable by assuming the curved edge lies on the $x=z$ plane. In this light, Theorem 4.7 used linear algebra to rearrange two trilinear equations to get Equations 77 and 74. As we mentioned in the manuscript, the new vertices lie on at least two hypersurfaces, and the new edges exist on the same hypersurface, while the eikonal constraint minimizes any associated error. This error is tolerated by the hyperparameter $\epsilon$=1e-4 (Section 6) as specified in Definition 3.4.
Figure 3 also suggests empirical evidence. Empirically, the flatness error, derived from the proof of Theorem D.5, is effectively controlled by the learning rate of the eikonal loss. In the experiments, we used 0.01 to give our results, while many other neural implicit surface learning methods were used to exploit a higher learning rate of 0.1 (e.g., Yariv et al. [7]). Note that, in practice, the eikonal loss is only applied to near-surface using clamps (if the target position is obviously far from a surface, ignore it; it is wise since we are concerned with the surfaces, not wanting exact SDFs for all space) for effective learning, which is sufficient to our extraction method.
Moreover, the smaller grid size of the Large model (ref. Section 6) in Figure 3 views a curved surface in a smaller interval, and it tends to regard it as a plane. From the plotting of Small to Large, our CD (colored crosses) is decreasing to zero, consistently with better CE, confirming this speculation.
B.2 On the claim that any encoding using trilinear interpolation can be applied
In this work, we describe “any encoding using trilinear interpolation (that) we defined can be applied.” The definition is implicitly referred to Definition 4.1 (Trilinear interpolation), where the cubic corners are represented by ${\bm{H}}\in\mathbb{R}^{D\times F}$. Since we do not rely on a specific mapping function $f:\mathrm{x}\rightarrow({\bm{H}}_{0},{\bm{H}}_{1}\dots\bm{H}_{7})$, we may apply our extraction method to any encoding using trilinear interpolation that we defined. Specifically, InstantNGP [12] used a hashing function to map a corner point to the arbitrary entry of hash tables, allowing gracious collisions with online adaptivity (refer to their Sec. 3), while TensoRF [13] can be interpreted that it used a mapping function to the summation of the multiplications of two vectors from the vector-matrix factorization. Therefore, if we have the full ranks for $R_{1},R_{2},R_{3}$, in Eqn. 3 in Chen et al. [13], among the VM factorization, which can be equivalently represented by the HashGrid with a sufficient number of hash table sizes.
Then, the discussion should be directed toward their encoding characteristics, which impact our extraction method due to their online adaptivity (HashGrid) or low-rank factorization (TensoRF). When we train an SDF, the optimization to global minima is generally difficult to achieve. Instead, previous works on neural implicit surface learning methods resort to optimizing the SDF on near-surface since we are interested in modeling the zero-set surfaces, allowing some errors for the points that are far from the surfaces. Practically, this is done by defining a loss function using clamps and ignoring far-surface points. For this reason, the online adaptivity of HashGrid or low-rank factorization of TensoRF can exploit this nature: focused learning of the surfaces adaptively allocating their learnable parameters to the near-surface, which means there might be enough room for the regularization by the eikonal equation. The investigation of the efficiency of representing the implicit neural surfaces using various positional encodings is seemingly out of the scope of this work since our main interest lies in the theoretical exposition of the mesh extraction from it and setting up a feasible bridge between theory and practice.
B.3 Zero-trailing binarization function
$\mathrm{B}$ is zero-trailing binarization function, consisting of $D=3$ bits. The zero-trailing means the number in bits is increasing left-aligned. For example, the number one is “100”, while the numbers two and four are “010” and “001”, respectively. The number three is “110”. From the left, each bit indicates x, y, and z axis in our notation. This convention seems to be widely used in the CUDA implementation in C language. Notice that we use the $j$ subscript of $\mathrm{B}_{j}$ to select the axis. Naturally, $i=0$ and $i=7$ indicate the opposite two corner points.
Appendix C Algorithms
Here are the algorithms described in Section 4.3 and Section 5.